# Kernel $k$-Medoids as General Vector Quantization

### v1

@@ I Introduction
Vector Quantization (VQ) [gersho2012vector, kohonen2001learning] is a foundational method in machine learning and data compression, where the goal is to represent a large dataset using a significantly smaller subset of representative points known as prototypes. In hard VQ, the prototypes are selected directly from the dataset itself, ensuring that all representations are grounded in real, observed data. These methods are particularly valued for their intuitive problem formulation and interpretable model behavior [lisboa2023coming]—traits that are increasingly important in domains such as healthcare [vellido2020importance], autonomous systems [kim2017interpretable] or scientific discovery [tshitoyan2019unsupervised], where transparency and explainability are essential. Depending on the context, prototypes may serve different purposes: in unsupervised learning, they are used for data representation or clustering [cheng1995mean], while in supervised settings, they support classification or regression tasks [macqueen1967some].
Among the many approaches to VQ, two families of methods have emerged as particularly prominent: distance-based centroid selection and probability density estimation methods. The former includes algorithms like $k$-means [lloyd1982least] and $k$-medoids [kaufman1987clustering], where the goal is to identify a set of data points that minimize the average dissimilarity to the rest of the dataset. The latter group includes Kernel Density Estimation (KDE) approaches, where the similarity between the data distribution and the selected prototypes is optimized using tools like the kernel trick and information-theoretic concepts [rao2007information, xu2008reproducing]. While both $k$-medoids-based VQ (MED-VQ) and KDE-based VQ (KDE-VQ) can be interpreted as prototype selection problems, their conceptual and mathematical connections remain poorly understood. They appear, at first glance, to address distinct optimization goals—centrality versus distribution matching—and have evolved largely independently within the literature.
A significant challenge shared by hard VQ methods is their combinatorial nature. Finding optimal prototype sets is computationally intractable in general, with problems like $k$-medoids being NP-hard [aloise2009np]. As the search space grows exponentially with the dataset size, exact solutions become impractical for real-world applications. This computational bottleneck has motivated the use of specialized hardware accelerators, particularly in solving Quadratic Unconstrained Binary Optimization (QUBO) [punnen2022quadratic] problems that arise in prototype selection, which are of the form
@@
where $\boldsymbol{Q}$ is an $n\times n$ real matrix. Among these hardware accelerators, Ising machines [mohseni2022ising] represent a notable direction. These devices are designed to solve optimization problems of the form given in Eq. 1 by finding the ground state of an Ising Hamiltonian, effectively translating optimization problems into physical energy minimization processes. Quantum Computing (QC) [nielsen2010quantum], in particular, offers a powerful model in this realm by exploiting superposition, entanglement, and quantum tunneling. One can differentiate between digital QC—coined Quantum Gate Computing (QGC)—which is used to implement sequential algorithmic structures via unitary operations and analog QC, such as Quantum Annealing (QA) [kadowaki1998quantum], based on the principles of Adiabatic Quantum Computing (AQC) [albash2018adiabatic]. While fully scalable quantum computers remain in development, early devices—classified as Noisy Intermediate-Scale Quantum (NISQ) hardware [preskill2018quantum]—are already accessible and have spurred growing interest in quantum optimization.
In this paper, we investigate the relationship between two QUBO formulations of hard VQ: one derived from $k$-medoids [bauckhage2019] clustering, and another from KDE divergence minimization [bauckhage2020]. The QUBO formulation for MED-VQ follows rather heuristic arguments where two competing objectives—the selection of central and mutually distant data points—are balanced via scalar weighting parameters. In opposition, a KDE-VQ QUBO is derived with a principled approach by minimizing the Maximum Mean Discrepancy (MMD) between the full dataset and a candidate prototype subset. Both VQ methods are first formulated as Quadratic Binary Programming (QBP) [laughhunn1970] problems and then reformulated to QUBOs by incorporating linear constraints. Surprisingly, we find that the principled QUBO based on MMD minimization is a special case of the heuristic MED-VQ approach. This equivalence emerges under a mild assumption: that the kernel-induced feature map embeds data onto the unit sphere of the feature space—an assumption that holds for commonly used Radial Basis Function (RBF) or Laplacian kernels as well as quantum kernels [havlivcek2019supervised]. Our contributions are threefold:
@@ II Related Work
A growing body of research explores how QC and quantum-inspired models can enhance or accelerate VQ. In [villmann2021quantum], the authors integrate VQ into a hybrid framework that leverages quantum-inspired models to optimize prototype adaptation. In [engelsberger2022steps], the authors take a step further by directly formulating learning VQ (LVQ) for classification tasks on theoretical QGC architectures. The work in [villmann2020quantum] explores similar ideas but focuses on how quantum principles, such as unitary operations and kernel-based feature embeddings, can be used to improve prototype-based classifiers.
Another promising direction involves leveraging AQC and related QUBO-based formulations to address the combinatorial nature of prototype selection in VQ. The foundation stone was layed in [bauckhage2019], where the authors come up with a QUBO formulation for the $k$-medoids problem, by balancing two competing objectives. This work is extended in [sijpesteijn2023quantum], where a preferable choice of the balancing parameters is proposed. A different approach is taken in [engelsberger2023quantum], where VQ is interpreted as a Set Cover problem, which is then reformulated into the QUBO framework. Finally, the authors of [bauckhage2020] utilize MMD for KDE to obtain a formulation suited for Hopfield networks, which are equivalent to QUBOs.
@@ III Background
[…]
@@ III-A $k$-Medoids
A local minimum of the objective in Eq. 2 can be obtained by using a slightly different versions of Lloyd’s algorithm [lloyd1982least]. This method initializes $k$ medoids and iteratively determines clusters by assigning data points to their closest medoid and updates these medoids according to that assignment. Finding a global optimum, however, is much harder, since $k$-medoids is NP-hard [aloise2009np].
The authors of [bauckhage2019] propose a QUBO formulation for estimating Eq. 2. For this, Fisher’s analysis of variance [fisher1921probable] is used, which states that the sum of the within cluster scatter and the between cluster scatter is constant. Minimizing the within cluster scatter can be written as selecting the $k$ most central data points
@@
[…]
@@
A QUBO formulation is obtained [bauckhage2019] by integrating the quadratic penalty function $(\boldsymbol{1}^{\top}\boldsymbol{z}-k)^{2}$ into the objective
@@
where equivalence to Eq. 4 holds for large enough $\lambda_{\text{MED}}>0$. Finding a suitable value is far from trivial and is of great interest in current research [alessandroni2023alleviating].
@@ III-B Kernel Density Estimation
As VQ aims to model the underlying probability density functions of the data distribution with the help of the codebook, we also consider probability density estimates. That is, assuming that $p_{\mathcal{D}}(\cdot)$ is an underlying probability density function of $\mathcal{D}$ and $p_{\mathcal{W}}(\cdot)$ of $\mathcal{W}$, we want $p_{\mathcal{D}}(\cdot)$ and $p_{\mathcal{W}}(\cdot)$ to be as similar as possible. The dissimilarity or divergence can be measured in numerous ways, examples include the Cauchy-Schwartz divergence [jenssen2006cauchy] or the Kullback-Leibler divergence [kullback1951information].
Since we have no knowledge on the underlying distribution of our data, we use approximations. In particular, we examine Kernel Density Estimation (KDE), also called Parzen windowing, similar to [bauckhage2020]
@@
where $K:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ is a kernel function. In [lehn2005vector, xu2008reproducing], an RBF kernel is used to approximate $p_{\mathcal{D}}(\cdot)$ and $p_{\mathcal{W}}(\cdot)$ and the Cauchy-Schwartz divergence is minimized. However, we are not restricted in the choice of the used kernel. Due to Mercer’s theorem [mercer1909xvi], there exists a feature map $\phi:\mathbb{X}\to\mathbb{R}^{D}$, s.t.,
@@
[…]
The mean vectors $\boldsymbol{\phi}_{\mathcal{X}}$ and $\boldsymbol{\phi}_{\mathcal{W}}$ fully characterize the KDEs $p_{\mathcal{D}}(\cdot)$ and $p_{\mathcal{W}}(\cdot)$. The difference between these vectors can be measured by the MMD [gretton2012kernel], which leads to the following objective
@@ IV Relationship between MED and KDE
[…]
@@
Using Theorem 1 and corollary 1, we can see that KDE-VQ is a special case of MED-VQ by choosing parameters in a specific way and using distance measures of the form $D(\boldsymbol{x},\boldsymbol{y})=1-K(\boldsymbol{x},\boldsymbol{y})$ for some Mercer kernel. With the commonly used RBF kernel with a bandwidth of $2$, one exactly recovers Welsch’s $M$-estimator, which was used as a distance function in [bauckhage2019].

### v2

@@ I Introduction
Vector Quantization (VQ) [1, 2] is a foundational method in machine learning and data compression, where the goal is to represent a large dataset using a significantly smaller subset of representative points known as prototypes. In hard VQ, the prototypes are selected directly from the dataset itself, ensuring that all representations are grounded in real, observed data. These methods are particularly valued for their intuitive problem formulation and interpretable model behavior [3]—traits that are increasingly important in domains such as healthcare [4], autonomous systems [5] or scientific discovery [6], where transparency and explainability are essential. Depending on the context, prototypes may serve different purposes: in unsupervised learning, they are used for data representation or clustering [7], while in supervised settings, they support classification or regression tasks [8].
Among the many approaches to VQ, two families of methods have emerged as particularly prominent: distance-based centroid selection and probability density estimation methods. The former includes algorithms like $k$-means [9] and $k$-medoids [10], where the goal is to identify a set of data points that minimize the average dissimilarity to the rest of the dataset. The latter group includes Kernel Density Estimation (KDE) approaches, where the similarity between the data distribution and the selected prototypes is optimized using tools like the kernel trick and information-theoretic concepts [11, 12]. While both $k$-medoids-based VQ (MED-VQ) and KDE-based VQ (KDE-VQ) can be interpreted as prototype selection problems, their conceptual and mathematical connections remain poorly understood. They appear, at first glance, to address distinct optimization goals—centrality versus distribution matching—and have evolved largely independently within the literature.
A significant challenge shared by hard VQ methods is their combinatorial nature. Finding optimal prototype sets is computationally intractable in general, with problems like $k$-medoids being NP-hard [13]. As the search space grows exponentially with the dataset size, exact solutions become impractical for real-world applications. This computational bottleneck has motivated the use of specialized hardware accelerators, particularly in solving Quadratic Unconstrained Binary Optimization (QUBO) [14] problems that arise in prototype selection, which are of the form
@@
where $\boldsymbol{Q}$ is an $n\times n$ real matrix. Among these hardware accelerators, Ising machines [15] represent a notable direction. These devices are designed to solve optimization problems of the form given in Eq. 1 by finding the ground state of an Ising Hamiltonian, effectively translating optimization problems into physical energy minimization processes. Quantum Computing (QC) [16], in particular, offers a powerful model in this realm by exploiting superposition, entanglement, and quantum tunneling. One can differentiate between digital QC—coined Quantum Gate Computing (QGC)—which is used to implement sequential algorithmic structures via unitary operations and analog QC, such as Quantum Annealing (QA) [17], based on the principles of Adiabatic Quantum Computing (AQC) [18]. While fully scalable quantum computers remain in development, early devices—classified as Noisy Intermediate-Scale Quantum (NISQ) hardware [19]—are already accessible and have spurred growing interest in quantum optimization applied to a diverse range of real-world problems [20, 21, 22, 23, 24].
In this paper, we investigate the relationship between two QUBO formulations of hard VQ: one derived from $k$-medoids [25] clustering, and another from KDE divergence minimization [26]. The QUBO formulation for MED-VQ follows rather heuristic arguments where two competing objectives—the selection of central and mutually distant data points—are balanced via scalar weighting parameters. In opposition, a KDE-VQ QUBO is derived with a principled approach by minimizing the Maximum Mean Discrepancy (MMD) between the full dataset and a candidate prototype subset. Both VQ methods are first formulated as Quadratic Binary Programming (QBP) [27] problems and then reformulated to QUBOs by incorporating linear constraints. Surprisingly, we find that the principled QUBO based on MMD minimization is a special case of the heuristic MED-VQ approach. This equivalence emerges under a mild assumption: that the kernel-induced feature map embeds data onto the unit sphere of the feature space—an assumption that holds for commonly used Radial Basis Function (RBF) or Laplacian kernels as well as quantum kernels [28]. Our contributions are threefold:
@@ II Related Work
A growing body of research explores how QC and quantum-inspired models can enhance or accelerate VQ. In [29], the authors integrate VQ into a hybrid framework that leverages quantum-inspired models to optimize prototype adaptation. In [30], the authors take a step further by directly formulating learning VQ (LVQ) for classification tasks on theoretical QGC architectures. The work in [31] explores similar ideas but focuses on how quantum principles, such as unitary operations and kernel-based feature embeddings, can be used to improve prototype-based classifiers.
Another promising direction involves leveraging AQC and related QUBO-based formulations to address the combinatorial nature of prototype selection in VQ. The foundation stone was layed in [25], where the authors come up with a QUBO formulation for the $k$-medoids problem, by balancing two competing objectives. This work is extended in [32], where a preferable choice of the balancing parameters is proposed. A different approach is taken in [33], where VQ is interpreted as a Set Cover problem, which is then reformulated into the QUBO framework. Finally, the authors of [26] utilize MMD for KDE to obtain a formulation suited for Hopfield networks, which are equivalent to QUBOs.
@@ III Background
For notational convenience, we indicate vectors by bold lowercase and matrices by bold capitalized letters. Further, we define $\mathbb{B}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{0,1\}^{n}$ and denote the vector consisting of only ones by $\boldsymbol{1}$, where its dimension is induced by the context. Moreover, the function $\operatorname*{diag}[\cdot]$, which takes a vector as input, denotes a diagonal matrix with that vector as its diagonal.
@@ III-A $k$-Medoids
A local minimum of the objective in Eq. 2 can be obtained by using a slightly different versions of Lloyd’s algorithm [9]. This method initializes $k$ medoids and iteratively determines clusters by assigning data points to their closest medoid and updates these medoids according to that assignment. Finding a global optimum, however, is much harder, since $k$-medoids is NP-hard [13].
The authors of [25] propose a QUBO formulation for estimating Eq. 2. For this, Fisher’s analysis of variance [34] is used, which states that the sum of the within cluster scatter and the between cluster scatter is constant. Minimizing the within cluster scatter can be written as selecting the $k$ most central data points
@@
by setting $\gamma\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{\beta}{\alpha}$. The parameter $\gamma$ then weighs the objectives, with a small value preferring the far away objective, while larger values lead to more centralized data points.
@@
A QUBO formulation is obtained [25] by integrating the quadratic penalty function $(\boldsymbol{1}^{\top}\boldsymbol{z}-k)^{2}$ into the objective
@@
where equivalence to Eq. 4 holds for large enough $\lambda_{\text{MED}}>0$. Finding a suitable value is far from trivial and is of great interest in current research [35].
@@ III-B Kernel Density Estimation
As VQ aims to model the underlying probability density functions of the data distribution with the help of the codebook, we also consider probability density estimates. That is, assuming that $p_{\mathcal{D}}(\cdot)$ is an underlying probability density function of $\mathcal{D}$ and $p_{\mathcal{W}}(\cdot)$ of $\mathcal{W}$, we want $p_{\mathcal{D}}(\cdot)$ and $p_{\mathcal{W}}(\cdot)$ to be as similar as possible. The dissimilarity or divergence can be measured in numerous ways, examples include the Cauchy-Schwartz divergence [36] or the Kullback-Leibler divergence [37].
Since we have no knowledge on the underlying distribution of our data, we use approximations. In particular, we examine Kernel Density Estimation (KDE), also called Parzen windowing, similar to [26]
@@
where $K:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ is a kernel function. In [38, 12], an RBF kernel is used to approximate $p_{\mathcal{D}}(\cdot)$ and $p_{\mathcal{W}}(\cdot)$ and the Cauchy-Schwartz divergence is minimized. However, we are not restricted in the choice of the used kernel. Due to Mercer’s theorem [39], there exists a feature map $\phi:\mathbb{X}\to\mathbb{R}^{D}$, s.t.,
@@
$\boldsymbol{\phi}_{\mathcal{D}}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{1}{n}\sum_{\boldsymbol{x}\in\mathcal{D}}\phi(\boldsymbol{x}),\quad\boldsymbol{\phi}_{\mathcal{W}}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{1}{k}\sum_{\boldsymbol{x}\in\mathcal{W}}\phi(\boldsymbol{x})\;.$
The mean vectors $\boldsymbol{\phi}_{\mathcal{X}}$ and $\boldsymbol{\phi}_{\mathcal{W}}$ fully characterize the KDEs $p_{\mathcal{D}}(\cdot)$ and $p_{\mathcal{W}}(\cdot)$. The difference between these vectors can be measured by the MMD [40], which leads to the following objective
@@ IV Relationship between MED and KDE
Let $K:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ be a normalized Mercer kernel, $\mathcal{D}=\{\boldsymbol{x}^{1},\dots,\boldsymbol{x}^{n}\}\subset\mathbb{\mathbb{X}}$, $k<n$ and $\boldsymbol{K}$ be the corresponding kernel matrix. . Defining the matrix $\boldsymbol{D}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{K}$, the QUBO matrices in Eq. 6 and Eq. 11 are equal if we set $\gamma=\frac{2k}{n}$ and $\lambda_{\text{KDE}}=\lambda_{\text{MED}}+1$.
@@
Using Theorem 1 and corollary 1, we can see that KDE-VQ is a special case of MED-VQ by choosing parameters in a specific way and using distance measures of the form $D(\boldsymbol{x},\boldsymbol{y})=1-K(\boldsymbol{x},\boldsymbol{y})$ for some Mercer kernel. With the commonly used RBF kernel with a bandwidth of $2$, one exactly recovers Welsch’s $M$-estimator, which was used as a distance function in [25].